OpenStudy (beth12345):

A boy and girl both swim at 3.0m/s. They jumpinto a river 1.0km across, with a current of 2.0m/s (E). She swims straight across relative to the shore. How long (in minutes) does each take to cross the river

OpenStudy (anonymous):

You are given the velocity of boy and girl w.r.t the ground and velocity of current w.r.t ground. Velocity of girl w.r.t ground, V(g,G) = 3 m/s Velocity of current = V (c,G) = 2 m/s Now, velocity of girl w.r.t current = V(g,G) + V(c,G) [ vector sum!! ] Hope you may try to solve it now...

OpenStudy (beth12345):

is it 5m per second

OpenStudy (anonymous):

No... I said the vector sum, as the direction of speeds are different you can't just add both the velocities..

OpenStudy (beth12345):

oh would I use a squared +b squared = c squared

OpenStudy (anonymous):

Exactly....

OpenStudy (beth12345):

i got 3.6

OpenStudy (anonymous):

But, think.. For calculating time she required to cross the river if she propells herself perpendicular to the shore, would in anyway be affected by velocity of current.

OpenStudy (anonymous):

Think why time depends on both components...

OpenStudy (beth12345):

uhm, I don`t understand

OpenStudy (beth12345):

 this is all I understand so far...